# Explicit quadratic large sieve inequality

### v1

@@ 1.1. Motivations and applications
and when $\mathcal{S}$ denotes all primitive characters with conductor $\leq Q$, there is
@@
Compared to the large sieve, the inequality (2) achieved a power saving in $Q$-aspect by relaxing the $\ell^{2}$-norm $\|a\|$. When $\{a_{n}\}$ is supported on squarefree integers, $\sum_{n_{1}n_{2}=\square}|a_{n_{1}}a_{n_{2}}|$ coincides with $\|a\|^{2}$. For general $\{a_{n}\}$, the quantity can be conveniently estimated. For instance, according to [4, p. 267], there is
@@
but $\|\mathbbm{1}_{\square}\|^{2}\leq\sqrt{N}$, so $A(S(Q),N,a)\ll_{\varepsilon}(QN)^{\varepsilon}(Q+N)\|a\|^{2}$ is generally false.
@@
If $f$ is an $SL_{2}(\mathbb{Z})$-cusp form of weight $\kappa$ that is also an eigenfunction to all Hecke operators with Fourier expansion
$f(z)=\sum_{n\geq 1}\lambda_{f}(n)n^{\kappa-1\over 2}e(nz)$
and $L(s,f\otimes\chi)$ be its $L$-function twisted by $\chi$:
@@
then an argument similar to the proof of (4) leads to
@@
In 2010, Soundararajan and Young [13] indicated that under GRH, when $\kappa\equiv 0\pmod{4}$ and $S^{*}(Q)$ denotes the collection of $\chi(\cdot)=\left(2\over\cdot\right)\left(\cdot\over m\right)$ with $1\leq m\leq Q$ squarefree odd, there is an asymptotic formula:
@@
The quadratic large sieve inequality (2) has recently gained renewed interest due to the breakthrough of Xiannan Li [7]. Specifically, by improving (2) for special choices of $a=\{a_{n}\}$, Li dropped the requirement of GRH for the asymptotic formula (6), which consequently turns many previously conditional results unconditional.
@@
For a prime $q\equiv 3\pmod{4}$, denote by $P_{q}(X)$ the number of rational prime $p\leq X$ that splits in $\mathbb{Q}(\sqrt{-q})$. Then Dunn et al. [2] used the quadratic large sieve (2) to show that for all $\varepsilon>0$, the asymptotic
@@ 1.2. Statement of results
Then there exists $C>0$ such that for every $\varepsilon>0$, we have
(11) $\mathcal{B}(M,N)\leq\exp_{4}(C\varepsilon^{-1})\cdot(MN)^{\varepsilon}(M+N).$
This improvement exhibits the explicit $\varepsilon$-dependence in the implied constants, and setting $\varepsilon=C_{1}/\log_{4}X$ for some large $C_{1}>0$ allows us to replace each $X^{\varepsilon}$-term in the estimates (2), (4), (7), (8), and (9) with $\exp(C_{2}\log X/\log_{4}X)$ for some $C_{2}>C_{1}$ when $X\geq\exp_{4}(1)$. To demonstrate this refinement, we prove the following:
@@
There exists $C>0$ such that for all $Q,N\geq 1$ satisfying $QN\geq\exp_{4}(1)$,
$\sum_{\chi\in S(Q)}\left|\sum_{n\leq N}a_{n}\chi(n)\right|^{2}\leq\exp_{4}(C(QN)/\log_{4}(QN))(Q+N)\sum_{\begin{subarray}{c}n_{1},n_{2}\leq N\\ n_{1}n_{2}=\square\end{subarray}}|a_{n_{1}}a_{n_{2}}|.$
@@
There exists $C>0$ such that when $s=\frac{1}{2}+it$, $T=|t|+1$, and $Q\geq\exp_{4}(1)$,
@@
There exists $C>0$ such that the asymptotic (9) is valid for all but $\ll Q\exp(-C\log Q/\log_{4}Q)$ amount of primes $q\in(Q,2Q]$ that are $\equiv 3\pmod{4}$, provided that $Q\to+\infty$ and $X\geq\exp(3C\log Q/\log_{4}Q)$.
@@ 1.3. Organization of the paper
In Section 2, Section 3, and Section 4, we present outline, lemmas, and the proof for Theorem 1. Then in Section 5, we derive the corollaries.
@@ 1.4. Acknowledgements
The work is supported by the LMS Undergraduate Research Bursary URB-2024-41. The author would like to express his sincere gratitude towards Professor Ian Petrow for suggesting the problem and many helpful comments.
@@ 1.5. Notations
  • $\|a\|=\left(\sum_{n}|a_{n}|^{2}\right)^{1/2}$ —– the $\ell^{2}$-norm for sequence $a=\{a_{n}\}$.
@@ 2.1. Preparation for Poisson summation
[…]
@@
Then there exists absolute $C>0$ such that for all $\varepsilon>0$, $K\leq M/2$, and $\Delta_{0}\leq N$, there is some $N_{1}\leq N/\Delta_{0}$ such that
@@ 2.2. Asymptotic developments
When $1\ll\Delta\ll N$ and $20<K\leq M/$ 2, there is
@@ 2.3. Recursive estimates
×$\times\Delta^{4}(M+N+M^{\frac{1}{2}}K^{\xi-\frac{1}{2}}+M^{\frac{1}{2}}NK^{-\frac{1}{2}}).$
@@
There exists $C_{2}>C_{1}>0$ such that for all $\varepsilon>0$ there is
@@ 3. Auxiliary lemmas
(21) $S(M,N,D,a,b)=\sum_{d\sim D}\sideset{}{{}^{*}}{\sum}_{m\sim M}\Biggl{|}\sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}(n_{1},n_{2})=1\\ d|n_{1}n_{2}\end{subarray}}a_{n_{1}}b_{n_{2}}\left(m\over n_{1}n_{2}\right)\Biggr{|}$
@@
$\leq\sum_{\delta\leq 2N}\sum_{\begin{subarray}{c}d_{1},d_{2}\\ d_{1}d_{2}\sim D\end{subarray}}\sideset{}{{}^{*}}{\sum}_{m\sim M}\Biggl{|}\sum_{\begin{subarray}{c}[\delta,d_{1}]|n_{1}\\ [\delta,d_{2}]|n_{2}\end{subarray}}a_{n_{1}}b_{n_{2}}\left(m\over n_{1}n_{2}\right)\Biggr{|}$
@@
By the relation (22), we see that each $\mu$ corresponds to a uniformly bounded range of $\nu$, so #$\#\{(\mu,\nu)\}=O(\log N)=O(\eta^{-1}N^{\eta/4})$, which means there exists some $P_{1},P_{2}$ such that
@@
$\Sigma_{a}=\sum_{\delta\leq 2N}\sum_{\begin{subarray}{c}d_{1}\sim P_{1}\\ d_{2}\sim P_{2}\end{subarray}}\sideset{}{{}^{*}}{\sum}_{m\sim M}\Biggl{|}\sum_{\begin{subarray}{c}n\\ [\delta,d_{1}]|n\end{subarray}}a_{n}\left(\frac{m}{n}\right)\Biggr{|}^{2}$
and $\Sigma_{b}$ is defined similarly. By the duality principle for bilinear forms (see [5, p. 170]) and Lemma 1, there is
@@
$=P_{2}\mathcal{B}(M,\theta N/P_{1})\sum_{n\sim N}|a_{n}|^{2}\sum_{\sigma|n}\mathop{\sum_{\delta\leq 2N}\sum_{d_{1}\sim P_{1}}}_{[\delta,d_{1}]=\sigma}1.$
@@
Finally, notice that when $n\sim N$ is squarefree, $4^{\omega(n)}=[d(n)]^{2}\leq\exp_{2}(C_{1}\eta^{-1})N^{\eta/2}$, so
@@
Let $W$ be a smooth function whose support is a compact subset of $(0,+\infty)$. Then for any positive squarefree odd $q$, there is
@@
If $A\geq 0$, $0<X_{1}\leq X\leq X_{2}$, $L\geq(X_{2}/X)^{2}$ and $J=\min(X/X_{1},X/X_{2})$, then for any integer $k\geq 2$, there is
@@
${\mathrm{d}^{n}\over\mathrm{d}x^{n}}[f(g(x))]=\sum_{\begin{subarray}{c}r_{1},r_{2},\dots,r_{n}\geq 0\\ r_{1}+2r_{2}+\cdots+nr_{n}=n\end{subarray}}{n!\over r_{1}!r_{2}!\cdots r_{n}!}f^{(n)}(g(x))\prod_{j=1}^{n}\left[g^{(j)}(x)\over j!\right]^{r_{j}}.$
In particular, when $g(x)$ is a quadratic polynomial, there is
${\mathrm{d}^{n}\over\mathrm{d}x^{n}}[f(g(x))]=\sum_{0\leq r\leq n/2}{n!\over 2^{r}r!(n-2r)!}f^{(n-r)}(g(x))[g^{\prime}(x)]^{n-2r}[g^{\prime\prime}(x)]^{r}.$
@@
For the general case, see [10, Theorem 2]. The following is a proof for the case where $g(x)$ is a quadratic polynomial:
@@
${\mathrm{d}\over\mathrm{d}x}$$\left\{f^{(n-r)}(g(x))[g^{\prime}(x)]^{n-2r}[g^{\prime\prime}(x)]^{r}\right\}$
@@
(25) ${\mathrm{d}^{n+1}\over\mathrm{d}x^{n+1}}[f(g(x))]$ $=f^{(n+1)}(g(x))[g^{\prime}(x)]^{n+1}$
@@
$\sum_{r=1}^{\mu}\left[{n!\over 2^{r}r!(n-2r)!}+{n!\over 2^{r-1}(r-1)!(n+1-2r)!}\right]$
×$\times f^{(n+1-r)}(g(x))[g^{\prime}(x)]^{n+1-2r}[g^{\prime\prime}(x)]^{r}$
@@
$=\sum_{0\leq r\leq(n+1)/2}{(n+1)!\over 2^{r}r!(n+1-2r)!}f^{(n+1-r)}(g(x))[g^{\prime}(x)]^{n+1-2r}[g^{\prime\prime}(x)]^{r},$
@@
so there is $W(x)=e^{\frac{8}{3}}\omega\left(x-\frac{3}{2}\right)$. Since for $10\leq x<$ 1,
@@
By Proposition 2, we deduce that for $10\leq x<$ 1,
@@
By Stirling’s approximation, when $|\eta|\ll|\xi|$, there is
@@ 4. Proof of Theorem 1
This section proves the results stated in Section 2. For the sake of bookkeeping, Section 4.X gives demonstrations of propositions in Section 2.X.
@@ 4.2. Asymptotic developments
in which $\sum_{a}$ sums over ±$\pm 1,\pm 2$ and
@@ 4.3. Recursive estimates
In this section, we prove Lemma 6, Lemma 7, and Lemma 8, which are needed to perform recursive estimates. We also derive Proposition 1 which analyzes the growth $F_{r}(\varepsilon)$ mentioned in Section 2.3.
@@
$\mathcal{B}(M,N_{0},K)\leq\exp_{2}(C_{3}\varepsilon^{-1})(MN_{0})^{C_{4}\varepsilon^{\frac{1}{2}}}F(\varepsilon)(M+M^{1-\xi}N_{0}^{2\xi-1}),$
@@
Therefore, in any case, there is
@@
which means when $r=\lfloor\varepsilon\rfloor$, there is
@@ 5.2. Moments of Dirichlet $L$-functions
In this section, we prove Corollary 2 by showing that when $s=\frac{1}{2}+it$, $T=|t|+1$, and $S^{\prime}(Q)=S(2Q)\smallsetminus S(Q)$ represents all primitive quadratic characters $\chi$ modulo $q_{\chi}\sim Q$, there is

### v2

@@ 1.1. Motivations and applications
and when $\mathcal{S}$ denotes all primitive characters with conductor $\leq Q$, we have
@@
Compared to the large sieve, the inequality (2) achieved a power saving in $Q$-aspect by relaxing the $\ell^{2}$-norm $\|a\|$. When $\{a_{n}\}$ is supported on squarefree integers, $\sum_{n_{1}n_{2}=\square}|a_{n_{1}}a_{n_{2}}|$ coincides with $\|a\|^{2}$. For general $\{a_{n}\}$, the quantity can be conveniently estimated. For instance, according to [4, p. 267], one has
@@
in which $p$ denotes a prime number. However, $\|\mathbbm{1}_{\square}\|^{2}\leq\sqrt{N}$, so $A(S(Q),N,a)\ll_{\varepsilon}(QN)^{\varepsilon}(Q+N)\|a\|^{2}$ is generally false.
@@
Let $f$ be an $SL_{2}(\mathbb{Z})$-cusp form of weight $\kappa$ that is also an eigenfunction to all Hecke operators with Fourier expansion
$f(z)=\sum_{n\geq 1}\lambda_{f}(n)n^{\kappa-1\over 2}e(nz),$
where $e(s)=e^{2\pi is}$, and let $L(s,f\otimes\chi)$ be its $L$-function twisted by $\chi$:
@@
where $e(s)=e^{2\pi is}$. Then an argument similar to the proof of (4) leads to
@@
In 2010, Soundararajan and Young [13] proved that under GRH, when $\kappa\equiv 0\pmod{4}$ and $S^{*}(Q)$ denotes the collection of $\chi(\cdot)=\left(2\over\cdot\right)\left(\cdot\over m\right)$ with $1\leq m\leq Q$ squarefree odd, one has an asymptotic formula:
@@
The quadratic large sieve inequality (2) has recently gained renewed interest due to the breakthrough of Li [7]. Specifically, by improving (2) for special choices of $a=\{a_{n}\}$, Li dropped the requirement of GRH for the asymptotic formula (6), which consequently turns many previously conditional results unconditional.
@@
For a prime $q\equiv 3\pmod{4}$, denote by $P_{q}(X)$ the number of rational primes $p\leq X$ that split in $\mathbb{Q}(\sqrt{-q})$. Then Dunn et al. [2] used the quadratic large sieve (2) to show that for all $\varepsilon>0$, the asymptotic
@@ 1.2. Statement of results
Then there exists a constant $C>0$ such that for every $\varepsilon>0$, we have
(11) $\mathcal{B}(M,N)\leq\exp_{4}(C\varepsilon^{-1})\cdot(MN)^{\varepsilon}(M+N),$
where $\exp_{k}$ refers to the k’th iterated exponential.
This improvement exhibits the explicit $\varepsilon$-dependence in the implied constants, and setting $\varepsilon=C_{1}/\log_{4}X$ ($\log_{k}$ refers to the k’th iterated logarithm) for some large constant $C_{1}>0$ allows us to replace each $X^{\varepsilon}$-term in the estimates (2), (4), (7), (8), and (9) with $\exp(C_{2}\log X/\log_{4}X)$ for some constant $C_{2}>C_{1}$ when $X\geq\exp_{4}(1)$. To demonstrate this refinement, we prove the following:
@@
There exists a constant $C>0$ such that for all $Q,N\geq 1$ satisfying $QN\geq\exp_{4}(1)$,
$\sum_{\chi\in S(Q)}\left|\sum_{n\leq N}a_{n}\chi(n)\right|^{2}\leq\exp(C(QN)/\log_{4}(QN))(Q+N)\sum_{\begin{subarray}{c}n_{1},n_{2}\leq N\\ n_{1}n_{2}=\square\end{subarray}}|a_{n_{1}}a_{n_{2}}|.$
@@
There exists a constant $C>0$ such that when $s=\frac{1}{2}+it$, $T=|t|+1$, and $Q\geq\exp_{4}(1)$,
@@
There exists a constant $C>0$ such that the asymptotic (9) is valid for all but $\ll Q\exp(-C\log Q/\log_{4}Q)$ amount of primes $q\in(Q,2Q]$ that are $\equiv 3\pmod{4}$, provided that $Q\to+\infty$ and $X\geq\exp(3C\log Q/\log_{4}Q)$.
@@ 1.3. Organization of the paper
In Section 2, Section 3, and Section 4, we present the outline, lemmas, and the proof for Theorem 1. Then in Section 5, we derive the corollaries.
@@ 1.4. Acknowledgements
The work is supported by the LMS Undergraduate Research Bursary URB-2024-41. The author would like to express his sincere gratitude towards Professor Ian Petrow for suggesting the problem and for many helpful discussions. The author also appreciates the anonymous reviewer for many useful comments.
@@ 1.5. Notations
  • $d(n)$ —– the divisor function $\sum_{d|n}1$.
  • $\omega(n)$ —– number of prime divisors of $n$. e.g. $\omega(2^{2}\times 3)=2$.
@@
  • $\|a\|=\left(\sum_{n}|a_{n}|^{2}\right)^{1/2}$ —– the $\ell^{2}$-norm for a sequence $a=\{a_{n}\}$.
@@ 2.1. Preparation for Poisson summation
$\Sigma_{1}$ $\leq\sideset{}{{}^{*}}{\sum}_{m\in\mathbb{Z}}W\left(\frac{m}{M}\right)\left|\sideset{}{{}^{*}}{\sum}_{n\sim N}a_{n}\left(\frac{m}{n}\right)\right|^{2}$
@@
Then there exists an absolute constant $C>0$ such that for all $\varepsilon>0$, $K\leq M/2$, and $\Delta_{0}\leq N$, there is some $N_{1}\leq N/\Delta_{0}$ such that
@@ 2.2. Asymptotic developments
When $1\ll\Delta\ll N$ and $0<K\leq M/2$, we have
@@ 2.3. Recursive estimates
$\times\Delta^{4}(M+N+M^{\frac{1}{2}}K^{\xi-\frac{1}{2}}+M^{\frac{1}{2}}NK^{-\frac{1}{2}}).$
@@
There exists $C_{2}>C_{1}>0$ such that for all $\varepsilon>0$ one has
@@ 3. Auxiliary lemmas
(21) $S(M,N,D,a,b)=\sum_{d\sim D}\sideset{}{{}^{*}}{\sum}_{m\sim M}\Biggl|\sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}(n_{1},n_{2})=1\\ d|n_{1}n_{2}\end{subarray}}a_{n_{1}}b_{n_{2}}\left(m\over n_{1}n_{2}\right)\Biggr|.$
@@
$\leq\sum_{\delta\leq 2N}\sum_{\begin{subarray}{c}d_{1},d_{2}\\ d_{1}d_{2}\sim D\end{subarray}}\sideset{}{{}^{*}}{\sum}_{m\sim M}\Biggl|\sum_{\begin{subarray}{c}[\delta,d_{1}]|n_{1}\\ [\delta,d_{2}]|n_{2}\end{subarray}}a_{n_{1}}b_{n_{2}}\left(m\over n_{1}n_{2}\right)\Biggr|.$
@@
By the relation (22), we see that each $\mu$ corresponds to a uniformly bounded range of $\nu$, so $\{(\mu,\nu)\}=O(\log N)=O(\eta^{-1}N^{\eta/4})$, which means there exists some $P_{1},P_{2}$ such that
@@
$\Sigma_{a}=\sum_{\delta\leq 2N}\sum_{\begin{subarray}{c}d_{1}\sim P_{1}\\ d_{2}\sim P_{2}\end{subarray}}\sideset{}{{}^{*}}{\sum}_{m\sim M}\Biggl|\sum_{\begin{subarray}{c}n\\ [\delta,d_{1}]|n\end{subarray}}a_{n}\left(\frac{m}{n}\right)\Biggr|^{2},$
and $\Sigma_{b}$ is defined similarly. By the duality principle for bilinear forms (see [5, p. 170]) and Lemma 1, we have
@@
$=P_{2}\mathcal{B}(M,\theta N/P_{1})\sum_{n\sim N}|a_{n}|^{2}\sum_{\sigma|n}\mathop{\sum_{\delta\leq 2N}\sum_{d_{1}\sim P_{1}}}_{[\delta,d_{1}]=\sigma}1$
@@
Finally, notice that when $n\sim N$ is squarefree, $4^{\omega(n)}=[d(n)]^{2}\leq\exp_{2}(C_{1}\eta^{-1})N^{\eta/2}$ [14, §5.2], so
@@
Let $W$ be a smooth function whose support is a compact subset of $(0,+\infty)$. Then for any positive squarefree odd $q$, we have
@@
If $A\geq 0$, $0<X_{1}\leq X\leq X_{2}$, $L\geq(X_{2}/X)^{2}$ and $J=\min(X/X_{1},X/X_{2})$, then for any integer $k\geq 2$, we have
@@
${\mathrm{d}^{n}\over\mathrm{d}x^{n}}(f\circ g)(x)=\sum_{\begin{subarray}{c}r_{1},r_{2},\dots,r_{n}\geq 0\\ r_{1}+2r_{2}+\cdots+nr_{n}=n\end{subarray}}{n!\over r_{1}!r_{2}!\cdots r_{n}!}f^{(n)}(g(x))\prod_{j=1}^{n}\left(g^{(j)}(x)\over j!\right)^{r_{j}}.$
In particular, when $g(x)$ is a quadratic polynomial, we have
${\mathrm{d}^{n}\over\mathrm{d}x^{n}}(f\circ g)(x)=\sum_{0\leq r\leq n/2}{n!\over 2^{r}r!(n-2r)!}f^{(n-r)}(g(x))[g^{\prime}(x)]^{n-2r}[g^{\prime\prime}(x)]^{r}.$
@@
For the general case, see [10, Theorem 2]. The following is a proof for the case where $g(x)$ is a quadratic polynomial.
@@
${\mathrm{d}\over\mathrm{d}x}$ $\left\{(f^{(n-r)}\circ g)(x)[g^{\prime}(x)]^{n-2r}[g^{\prime\prime}(x)]^{r}\right\}$
@@
(25) ${\mathrm{d}^{n+1}\over\mathrm{d}x^{n+1}}(f\circ g)(x)$ $=f^{(n+1)}(g(x))[g^{\prime}(x)]^{n+1}$
@@
$\sum_{r=1}^{\mu}\left({n!\over 2^{r}r!(n-2r)!}+{n!\over 2^{r-1}(r-1)!(n+1-2r)!}\right)$
$\times f^{(n+1-r)}(g(x))[g^{\prime}(x)]^{n+1-2r}[g^{\prime\prime}(x)]^{r}$
@@
$\sum_{0\leq r\leq(n+1)/2}{(n+1)!\over 2^{r}r!(n+1-2r)!}f^{(n+1-r)}(g(x))[g^{\prime}(x)]^{n+1-2r}[g^{\prime\prime}(x)]^{r},$
@@
so $W(x)=e^{\frac{8}{3}}\omega\left(x-\frac{3}{2}\right)$. Since for $0\leq x<1$,
@@
By Proposition 2, we deduce that for $0\leq x<1$,
@@
By Stirling’s approximation [15, p. 78], when $|\eta|\ll|\xi|$, we have
@@ 4. Proof of Theorem 1
This section proves the results stated in Section 2. For the sake of bookkeeping, Section 4.X gives demonstrations of lemmas and Proposition 1 in Section 2.X.
@@ 4.2. Asymptotic developments
in which $\sum_{a}$ sums over $\pm 1,\pm 2$ and
@@ 4.3. Recursive estimates
In this section, we prove Lemma 6, Lemma 7, and Lemma 8, which are needed to perform recursive estimates. We also derive Proposition 1, which analyzes the growth $F_{r}(\varepsilon)$ mentioned in Section 2.3.
@@
$\mathcal{B}(M,N_{0},K)\leq\exp_{2}(C_{3}\varepsilon^{-1})(MN_{0})^{C_{4}\varepsilon^{\frac{1}{2}}}F(\varepsilon)(M+M^{1-\xi}N_{0}^{2\xi-1}).$
@@
Therefore, in any case, we have
@@
which means when $r=\lfloor\varepsilon^{-1}\rfloor$, we have
@@ 5.2. Moments of Dirichlet $L$-functions
In this section, we prove Corollary 2 by showing that when $s=\frac{1}{2}+it$, $T=|t|+1$, and $S^{\prime}(Q)=S(2Q)\smallsetminus S(Q)$ represents all primitive quadratic characters $\chi$ modulo $q_{\chi}\sim Q$, one has